The quantity produced daily at the ABC cement factory is approximately normally



distributed with mean 0.82 and standard deviation 0.14. The units are on millions of tons.



Find the probability that the total production will between 0.80 and 0.85 million tons.

1
Expert's answer
2022-04-07T05:43:03-0400

We have a normal distribution, "\mu=0.82, \sigma=0.14."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma};"

"z_1=\cfrac{0.80-0.82}{0.14}=-0.14;\\
z_2=\cfrac{0.85-0.82}{0.14}=0.21;\\
P(0.80<X<0.85)=P(-0.14<Z<0.21)=\\
=P(Z<0.21)-P(Z<-0.14)="

"=0.5832-0.4443=0.1389" ​(from z-table).